# Surveying the space of descriptions of a composite system with machine learning

### v1

@@ Abstract
The codebase is written to expect the first $N$ terms to be the individual components, as these have a special role in the training loss for two reasons. First, the sum total information from all components is driven to a specified value $\hat{I}_{\text{in}}$, rather than extremized as is done for the remaining terms. Second, the component-wise information terms $I(U_{i};X_{i})$ are estimated more directly than the remaining terms, since we have access to the posterior distributions. For these, we use the lower bound in Section 2.5 of Poole et al. [26], while the remaining mutual information terms are estimated through InfoNCE [27].
The training process. For the spin system, a single optimization was run for all transmitted information values $\hat{I}_{\text{in}}$ for each extremization (minimization or maximization) of a given quantity. The value was increased linearly from zero bits to five bits over the course of training. For Sudoku and the $n$-gram statistics, we found it necessary to run separate optimizations for each value of $\hat{I}_{\text{in}}$, and to increase the coefficient $\gamma$ on the sum total transmitted information in stages over the course of training. During each optimization, $\gamma$ was held fixed at a low value $\gamma_{0}$ for the first half of training. Then it was increased exponentially to its final value $\gamma_{1}$ over the second half of training. After training, the latent codes (specifying the description) were fixed, and then the remaining neural networks continued to train with a larger batch size in order to refine the estimates of their corresponding mutual information terms. We found this necessary for Sudoku and $n$-grams, and particularly for O-information, where errors on the many mutual information terms can easily dominate the estimate.
Training hyperparameters and architecture details. For each of the three systems, we list implementation specifics in Tables 1-3. The “Encoder MLP architecture” was used for every MLP involved in the InfoNCE terms (two MLPs per term).
@@
Transmitted information coefficient $\gamma_{0}$ 0.5
@@
Description space in terms of total correlation, Sudoku and 4-grams. Figs. S1&S2 are the counterparts to Figs. 2&3 in the main text, where the description spaces for Sudoku and 4-grams are visualized and extremized in terms of total correlation instead of O-information.
For Sudoku (Fig. S1), the most redundant descriptions from before are also the ones with largest total correlation. However, the most synergistic descriptions found before (schemes v and vi) lie internal to the total correlation boundaries. The subsystem with minimal total correlation for eight bits of information (four squares), scheme vii, shows the most independent squares on the board.
@@
Regarding the 4-grams, there is less contrast between the 4-letter and 8-letter words when viewing their description spaces in terms of total correlation than there was in terms of O-information. We include an additional reference scheme in this space: the diamond in Fig. S2a represents a vowel/consonant compression scheme for all four letters, and is near the maximum total correlation found by the optimization. In Fig. S2b, we show four of the hardened compression schemes; note how individual letters are distinguished for maximal total correlation while multiple letters are grouped for pieces of information that have minimal total correlation.
Supplemental Material, Figure S2: Statistics of 4-letter sequences in terms of total correlation. (a) The space of descriptions for 4-grams taken from 4- and 8-letter words, plotted in terms of total correlation. Vowel/consonant refers to a description based on distinguishing vowels from consonants for each letter. (b) The hardened compression schemes for the minimal and maximal O-information points marked as stars in panel (a), where groupings of letters are separated by vertical bars, and the group with all remaining letters of the alphabet is represented by an asterisk (*).

### v2

@@ Abstract
The codebase is written to expect the first $N$ terms to be the individual components, as these have a special role in the training loss for two reasons. First, the sum total information from all components is driven to a specified value $\hat{I}_{\text{in}}$, rather than extremized as is done for the remaining terms. Second, the component-wise information terms $I(U_{i};X_{i})$ are estimated with the lower bound in Section 2.5 of Poole et al. [33], while the remaining mutual information terms are estimated through InfoNCE [34].
The training process. For the spin system, a single optimization was run for all transmitted information values $\hat{I}_{\text{in}}$ for each extremization (minimization or maximization) of a given quantity. The value was increased linearly from zero bits to five bits over the course of training. For sudoku and the $n$-gram statistics, we found it necessary to run separate optimizations for each value of $\hat{I}_{\text{in}}$, and to increase the coefficient $\gamma$ on the sum total transmitted information in stages over the course of training. During each optimization, $\gamma$ was held fixed at a low value $\gamma_{0}$ for the first half of training. Then it was increased exponentially to its final value $\gamma_{1}$ over the second half of training.
For evaluation, we sampled data points $x\sim p(x)$ and then embeddings $u\sim p(u|x)$ to compute the expectation
$I(X;U)=\mathbb{E}_{x,u\sim p(x,u)}[\log\frac{p(u|x)}{p(u)}],$ (6)
with $p(u)=\sum_{i}^{M}p(x_{i})p(u|x_{i})$ aggregated over the entire dataset. For all analyses, we sampled $2\times 10^{5}$ points and used the standard error of the estimate as its uncertainty. Error propagation then gave the uncertainties on the summary quantities and total component information.
For the spin systems and sudoku, we repeated each run five times and used the best performer, which was simply the point (or scan) that yielded the maximal/minimal summary quantity. For n-grams, we trained once, albeit after some hyperparameter tuning.
Training hyperparameters and architecture details. For each of the three systems, we list implementation specifics in Tables S1-3. The “Encoder MLP architecture” was used for every MLP involved in the InfoNCE terms (two MLPs per term).
@@
Transmitted information coefficient $\gamma_{0}$ 2
@@
Sudoku description space in terms of total correlation. Fig. S1 is the counterpart to Fig. 2 in the main text, where the description space for sudoku is visualized and extremized in terms of total correlation instead of O-information.
The most redundant descriptions from before are also the ones with largest total correlation. However, the most synergistic descriptions found before (schemes v and vi) lie internal to the total correlation boundaries. The subsystem with minimal total correlation for eight bits of information (four squares), scheme vii, shows the most independent squares on the board.
@@
Supplemental Material, Figure S2: Structure of 4-grams statistics, continued. We reproduce the total correlation (a) and O-information (b) description spaces from Fig. 3 in the main text, now with the discrete subsystems for all three datasets (squares), and with hardened descriptions for all extremized quantities at around seven bits of total information (stars). (c) For the 4-letter words, we hardened the descriptions that minimize and maximize total correlation and O-information, and show the top contributing codes.
Supplemental Material, Figure S3: Structure of 4-grams statistics, continued. For the first half (a) and second half (b) of 8-letter words, we hardened the descriptions that minimize and maximize total correlation and O-information, and show the top contributing codes.
Contribution of specific codes to summary quantities (Fig. 3, main text). By expressing total correlation $\text{TC}(\bm{U})$ as an expectation over codes $\bm{u}$, we can compute the contribution per code for additional insight about the variation that a description encapsulates [19, 43]. To be specific,
$\text{TC}(\bm{U})=\mathbb{E}_{\bm{u}\sim p(\bm{u})}[\log\frac{p(\bm{u})}{\Pi_{i}^{N}p(u_{i})}]=\mathbb{E}_{\bm{u}\sim p(\bm{u})}[\text{tc}(\bm{u})],$ (7)
framing total correlation as a comparison between the probability of outcome $\bm{u}$ when accounting for all components jointly, $p(\bm{u})$, versus independently, $\Pi_{i}^{N}p(u_{i})$. We then evaluate the contribution of each outcome to total correlation as $\Delta\text{TC}(\bm{u})=p(\bm{u})\cdot\text{tc}(\bm{u})$.
O-information permits a similar framing [43], now as a comparison between the joint and product-of-marginals probabilities, and the joint and marginalize-one-out probabilities:
$\Omega(\bm{U})=\mathbb{E}_{\bm{u}\sim p(\bm{u})}[2\log\frac{p(\bm{u})}{\Pi_{i}^{N}p(u_{i})}-\sum_{i}^{N}\log\frac{p(\bm{u})}{p(\bm{u}_{/i})p(u_{i})}]=\mathbb{E}_{\bm{u}\sim p(\bm{u})}[\omega(\bm{u})].$ (8)
We show the sorted contributions to total correlation ($\Delta$TC) and O-information ($\Delta\Omega$), and the groupings of 4-grams that contribute most on either end of the spectrum, in Figs. S2&S3.
S1 Citation Diversity Statement
Science is a human endeavour and consequently vulnerable to many forms of bias; the responsible scientist identifies and mitigates such bias wherever possible. Meta-analyses of research in multiple fields have measured significant bias in how research works are cited, to the detriment of scholars in minority groups [45, 46, 47, 48, 49, 50]. We use this space to amplify studies, perspectives, and tools that we found influential during the execution of this research [51, 52, 53, 54]. We sought to proactively consider choosing references that reflect the diversity of the field in thought, form of contribution, gender, race, ethnicity, and other factors. The gender balance of papers cited within this work was quantified using a combination of automated gender-api.com estimation and manual gender determination from authors’ publicly available pronouns. By this measure (and excluding self-citations to the first and last authors of our current paper), the references of the main text contain 5% woman(first)/woman(last), 10% man/woman, 18% woman/man, and 68% man/man. This method is limited in that a) names, pronouns, and social media profiles used to construct the databases may not, in every case, be indicative of gender identity and b) it cannot account for intersex, non-binary, or transgender people. We look forward to future work that could help us to better understand how to support equitable practices in science.